OpenStudy (anonymous):

@Mathematics Day 1 2 3 4 5 Temperature (°F) 70 72 68 75 70 The temperature for 5 days is shown in the table. If on the 6th day the temperature is 83°, how will this change the overall mean temperature? A)increase by 1° B)increase by 2° C)increase by 3° D)increase by 4°

OpenStudy (anonymous):

70 + 72 +68 +75 +70 = 355 355/5 = 71 70 + 72 +68 +75 +70 + 83 = 438 438/6 = 73 73 - 71 = 2 So the answer is B